calculate the pH of an aqueous solution of 0.1 M ammonium formate assuming complete dissociation.pkb for formic acid=3.8,and pkb of ammonia=4.8?

Grade:12

2 Answers

Sunil Kumar FP
10 years ago

wallt of weak acid and weak base
ph=1/2pkw + 1/2pka -1/2pkb
=1/2×14 + 1/2×3.8 -1/2×4.8
=6.5

thanks and regards
sunil kr
askIItian faculty
ankit singh
3 years ago

When it comes to Ammonium formate, it is the perfect blend of weak base and weak acid.  According to the pH formula, we need to substitute the values.  When we place the values, we can get the following equation.  The given conditions are the pka of formic acid = 3.8 and pkb of ammonia = 4.8

ph=1/2pkw + 1/2pka -1/2pkb

=1/2×14 + 1/2×3.8 -1/2×4.8

pH=6.5.